Comparison of PRN Techniques for Small Size PPS Sample Coordination

Esbjörn Ohlsson^{∗}

March 1999

Abstract

Consider multi-stage sampling from a stratified finite population, with a few primary sampling units selected in each stratum using probabilities proportional to size (pps). In a repeated survey, it is at times desired to redesign the sample with new size measures and new strata, while retaining as many units as possible from the old sample. In a former paper (Ohlsson, 1996) we considered the case with sample size n = 1, for which we gave an overview of existing methods and proposed a new method based on the use of permanent random numbers (PRN). In the present paper we focus on the case with small stratum sample sizes (2 ≤ n ≤ 4). We discuss the properties of different PRN methods and present a simulation study of their achieved sample overlap.

KEY WORDS: overlap maximization; overlap control; sample redesign; permanent random numbers; probabilities proportional to size.

∗Address: Mathematical Statistics, Stockholm University, S-106 91 Stockholm, Sweden.

e-mail: esbj@matematik.su.se. Financial support from the Bank of Sweden Tercentenary Foundation is gratefully acknowledged.